# Theoretical Insights into Fine-Tuning Attention Mechanism: Generalization and Optimization

### v1

@@ Abstract
Large Language Models (LLMs), built on Transformer architectures, exhibit remarkable generalization across a wide range of tasks. However, fine-tuning these models for specific tasks remains resource-intensive due to their extensive parameterization. In this paper, we investigate two remarkable phenomena observed during the fine-tuning of LLMs, particularly focusing on the attention mechanism: (1) Different Impact, optimizing the $\mathbf{W}_{v}$ matrix significantly improves performance over optimizing the $\mathbf{W}_{k}$ matrix. Fine-tuning only the $\mathbf{W}_{q}$ and $\mathbf{W}_{v}$ matrices is computationally efficient, delivering results that are comparable to, or even better than, fine-tuning all three matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$. (2) Efficient Convergence, employing distinct learning rates for these matrices is crucial for optimal performance, with a higher learning rate for the $\mathbf{W}_{v}$ matrix expediting convergence. However, theoretical analyses of these phenomena are still relatively limited. We present a theoretical analysis of these phenomena from two perspectives: (i) Generalization, where we demonstrate that fine-tuning only $\mathbf{W}_{q}$ and $\mathbf{W}_{v}$ improves generalization bounds, enhances memory efficiency, and (ii) Optimization, where we emphasize that the feature learning of the attention mechanism is efficient, particularly when using distinct learning rates for the matrices, which leads to more effective fine-tuning. Building on these insights, we propose a new strategy that improves fine-tuning efficiency in terms of both storage and time. Experimental results on benchmark datasets validate the effectiveness of this approach, supporting our theoretical findings. Our analysis lays the theoretical groundwork for configuring and improving lightweight algorithms in LLMs fine-tuning.
@@ 1 Introduction
A fundamental component of transformers is the attention mechanism, particularly the interactions among the query matrix $\mathbf{W}_{q}$, the key matrix $\mathbf{W}_{k}$, and the value matrix $\mathbf{W}_{v}$. During the fine-tuning of LLMs involving the attention mechanism, two interesting phenomena have been observed: (1) Different Impact—optimizing the $\mathbf{W}_{v}$ matrix is pivotal for enhancing performance, significantly more so than adjustments to the $\mathbf{W}_{k}$ matrix, which exhibit limited impact on the outcomes. Additionally, fine-tuning only the $\mathbf{W}_{q}$ and $\mathbf{W}_{v}$ matrices often yields results that are comparable to or surpass those achieved by fine-tuning all three matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$, which also reduces the number of tunable parameters by approximately 1/3, offering computational benefits (Section 3). (2) Efficient Convergence—using the same learning rate for $\mathbf{W}_{q}\&\mathbf{W}_{k}$ and $\mathbf{W}_{v}$ is not optimal for efficient convergence. In fact, it is essential to apply distinct learning rates for the $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$ components to ensure optimal fine-tuning performance. Specifically, the learning rate for $\mathbf{W}_{v}$ should generally be higher than that for $\mathbf{W}_{q}$ and $\mathbf{W}_{k}$ to facilitate efficient convergence (Section 4).
@@
While certain empirical guidelines, such as the original Low-Rank Adaptation (LoRA) [18], explore which weight matrices in transformers are suitable for the application of LoRA, comprehensive theoretical analyses of these phenomena are still limited. This includes aspects such as selecting appropriate weight types for fine-tuning and optimizing learning rate settings. In this paper, we seek to explore the issue from the following two aspects.
@@
  • Generalization: advantages of fine-tuning $\mathbf{W}_{q}\&\mathbf{W}_{v}$ over $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$ together. We perform a thorough theoretical analysis to demonstrate the advantages. To be more specific, we employ information-theoretic approaches [42, 31, 41, 48] to establish the generalization bounds of fine-tuning pre-trained models with attention mechanism (See Theorem 1 for details). This indicates that fine-tuning $\mathbf{W}_{q}\&\mathbf{W}_{v}$ instead of $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$ reduces the number of parameters, while improving generalization bounds and potentially providing memory benefits.
@@
  • Optimization: convergence analysis of attention mechanism with varying learning rate settings. To further investigate the aforementioned phenomena, we examine the optimization process of the attention mechanism. First, we discuss the learning dynamics in transformers in Case 1, suggest that $\mathbf{W}_{v}$ may experience instances of inefficient learning during fine-tuning process for downstream tasks. This naturally leads to the hypothesis that accelerating the learning of $\mathbf{W}_{v}$ in the early stages could potentially induce $\mathbf{W}_{k}$ and $\mathbf{W}_{q}$ to begin learning earlier. Additionally, by using scaling arguments for large width-$n$ networks [45, 11], we illustrate (Theorem 2) that the feature learning of attention mechanism is efficient when the learning rate for $\mathbf{W}_{v}$ should be generally much larger than that of $\mathbf{W}_{q}\&\mathbf{W}_{k}$ in fine-tuning.
@@
Building on our experimental and theoretical insights, one can develop new algorithms to improve the effectiveness (e.g., storage, and time) of fine-tuning. Experimental results for our strategy (in Section 5) on benchmark datasets [40] and open source pre-trained models [26, 2] verify that the method can visibly influence fine-tuning efficiency [^1] The code is available at https://github.com/chen123CtrlS/LightweightAtt. We do not make direct comparisons with various parameter-efficient fine-tuning methods, as our strategy is primarily intended to demonstrate how theoretical analysis can effectively guide experimental procedures.
@@ 2 Preliminaries and Background
Self-attention. Given a sequence of $m$ vectors $\mathbf{C}\in\mathbb{R}^{m\times d_{in}}$ over which we would like to perform attention and a query vector $\mathbf{x}\in\mathbb{R}^{d_{in}}$, that is, the input is [$\mathbf{C},\mathbf{x}$] $\in\mathbb{R}^{(m+1)\times d_{in}}$. The conventional attention function can be expressed as [^2] For simplicity, we focus on the last vector of input in a single-head self-attention. Our analysis is readily generalizable to multi-head self-attention.:
$\text{Attn}(\mathbf{x}\mathbf{W}_{q},\mathbf{C}\mathbf{W}_{k},\mathbf{C}\mathbf{W}_{v})\\ =\text{softmax}\left(\frac{\mathbf{x}\mathbf{W}_{q}\mathbf{W}_{k}^{T}\mathbf{C}^{T}}{\sqrt{d_{out}}}\right)\mathbf{CW}_{v},$ (1)
@@
A unified framework for parameter-efficient fine-tuning. Building on the work of [14], we consider a unified framework that establishes connections among various parameter-efficient fine-tuning methods. Specifically, we reinterpret these methods as modifications applied to specific hidden states within pre-trained models, the composition function:
@@
where $l_{1},l_{2}$ are coefficients, $\mathbf{h}$ is denoted as the hidden representation to be directly modified and $\Delta\mathbf{h}$ is a modification vector. Additionally, $\mathbf{h}$ and $\mathbf{x}$ can represent the attention output and input respectively. Here, we will present two special cases:
@@
    $\mathbf{h}\leftarrow(1-\lambda(\mathbf{x})\mathbf{h}+\lambda(\mathbf{x})\Delta\mathbf{h},\ \ \ \Delta\mathbf{h}:=\text{softmax}(\mathbf{xW}_{q}\mathbf{P}_{k}^{T})\mathbf{P}_{v}\triangleq\text{softmax}(\mathbf{xA})\mathbf{B},$ (4)
    where $\lambda(\mathbf{x})=\frac{\sum_{i}\text{exp}(\mathbf{xW}_{q}\mathbf{P}_{k}^{T})_{i}}{\sum_{i}\text{exp}(\mathbf{xW}_{q}\mathbf{P}_{k}^{T})_{i}+\sum_{j}\text{exp}(\mathbf{xW}_{q}\mathbf{W}_{k}^{T}\mathbf{C}^{T})_{j}}$ is a scalar that represents the sum of normalized attention weights on the prefixes. We derive a detailed equivalent form of Prefix tuning to establish its connection with LoRA in Appendix B.1.
@@
By defining $\mathbf{A}=\mathbf{W}_{q}\mathbf{P}_{k}^{T},\mathbf{B}=\mathbf{P}_{v}$ in Eq.(4), we can establish a connection with LoRA in Eq.(3). Notably, if we replace the softmax attention with linear attention here, the two are equivalent to some extent. Intuitively, in the attention mechanism, $\mathbf{A}\ (\mathbf{W}_{q}\mathbf{P}_{k}^{T})$ is responsible for generating attention scores, while $\mathbf{B}\ (\mathbf{P}_{v})$ utilizes these attention scores to produce the target content. Therefore, during fine-tuning, query, key, and value are likely to exhibit varying degrees of importance. This may also provide theoretical insights for recent works [48, 10], which empirically observed an asymmetry where the project-down matrix $\mathbf{A}$ is responsible for extracting features from the input, while the project-up matrix $\mathbf{B}$ utilizes these features to generate the desired output in LoRA fine-tuning.
$\Theta$ Notation. For the convergence analysis, we adopt the following notation to describe the asymptotic behavior as the width $n$ increases, similar to those in [45, 11]. Given sequences $c_{n}\in\mathbb{R}$ and $d_{n}\in\mathbb{R}^{+}$, we write $c_{n}=O(d_{n})$ and $c_{n}=\Omega(d_{n})$ to mean $c_{n}<\kappa d_{n}$ or $c_{n}>\kappa d_{n}$, respectively, for some constant $\kappa>0$. We denote $c_{n}=\Theta(d_{n})$ when both $c_{n}=O(d_{n})$ and $c_{n}=\Omega(d_{n})$ hold, implying that $c_{n}$ and $d_{n}$ grow at comparable rates. For vector sequences $c_{n}=(c_{n}^{i})_{1\leq i\leq k}\in\mathbb{R}^{k}$ (for some $k>0$), we write $c_{n}=O(d_{n})$ when $c_{n}^{i}=O(d_{n}^{i})$ for all $i\in[k]$, and analogous notation applies for other asymptotic bounds. Finally, when the sequence $c_{n}$ is a vector of random variables, convergence is understood to refer to convergence in the second moment (i.e., $L_{2}$ norm).
@@ 3 Advantages and Generalization Analysis
This section, we show our first interesting observation (Different Impact) in fine-tuning the attention mechanism and the storage benefit of fine-tuning only $\mathbf{W}_{q}$ and $\mathbf{W}_{v}$ in Section 3.1. Afterwards, we give a mutual information based generalization bounds of fine-tuning only $\mathbf{W}_{q}$ and $\mathbf{W}_{v}$ in Section 3.2, which provide a better generalization error.
3.1 Empirical Advantages of Fine-Tuning only Query,Value Matrices
To explore the Different Impact, we focus our study on adapting only the attention weights for downstream tasks, while freezing the other modules to ensure simplicity and parameter efficiency. Furthermore, we investigate the impact of adapting different types of attention weight matrices in a Transformer, as outlined below. We present our empirical results using LoRA to fine-tune a set of language models (Roberta-base [26] and Llama3.1-8b [2]) across various benchmarks [40]. Further details on the experimental setup and additional empirical results can be found in Appendix C.1.

### v3

@@ Abstract
Large Language Models (LLMs), built on Transformer architectures, exhibit remarkable generalization across a wide range of tasks. However, fine-tuning these models for specific tasks remains resource-intensive due to their extensive parameterization. In this paper, we explore two remarkable phenomena related to the attention mechanism during the fine-tuning of LLMs (where $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$ denote the weights of the query, key, and value layers, respectively). The first phenomenon, termed “Unequal Importance of Attention Matrices”, highlights the impact of fine-tuning different weight matrices. It shows that optimizing the $\mathbf{W}_{v}$ matrix yields significantly better performance than optimizing the $\mathbf{W}_{k}$ matrix. Fine-tuning only the $\mathbf{W}_{q}$ and $\mathbf{W}_{v}$ matrices is computationally efficient while delivering results comparable to, or even better than fine-tuning all three matrices ($\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$). The second phenomenon, “Attention Matrices with Customized Learning Rate Lead to Better Convergence”, emphasizes the importance of assigning distinct learning rates to these matrices. Specifically, a higher learning rate for the $\mathbf{W}_{v}$ matrix compared to $\mathbf{W}_{q}$ and $\mathbf{W}_{k}$ accelerates convergence and improves performance. Building on these insights, we propose a new strategy that improves fine-tuning efficiency in terms of both storage and time. Experimental results on benchmark datasets validate the effectiveness of this approach, supporting our theoretical findings. Our analysis lays the theoretical groundwork for configuring and improving algorithms in LLMs fine-tuning [^3] Extended version and code, are available at https://github.com/chen123CtrlS/EfficientFT/.
@@ 1 Introduction
A fundamental component of transformers is the attention mechanism, particularly the interactions among the query matrix $\mathbf{W}_{q}$, the key matrix $\mathbf{W}_{k}$, and the value matrix $\mathbf{W}_{v}$. During the fine-tuning of LLMs involving the attention mechanism, two interesting phenomena have been observed: (1) Unequal Importance of Attention Matrices—optimizing the $\mathbf{W}_{v}$ is pivotal for enhancing performance, significantly more so than adjustments to the $\mathbf{W}_{k}$, which exhibit limited impact on the outcomes. Additionally, fine-tuning only the $\mathbf{W}_{q}$ and $\mathbf{W}_{v}$ often yields results that are comparable to or surpass those achieved by fine-tuning all three matrices $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$, which also reduces the number of tunable attention parameters by approximately 1/3, offering computational benefits (Section 3). (2) Attention Matrices with Customized Learning Rate Lead to Better Convergence—using the same learning rate for $\mathbf{W}_{q}\&\mathbf{W}_{k}$ and $\mathbf{W}_{v}$ is not optimal for efficient convergence. In fact, it is essential to apply distinct learning rates for the $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$ components to ensure optimal fine-tuning performance. Specifically, the learning rate for $\mathbf{W}_{v}$ should generally be higher than that for $\mathbf{W}_{q}$ and $\mathbf{W}_{k}$ to facilitate efficient convergence (Section 4).
@@
While certain empirical guidelines, such as the original Low-Rank Adaptation (LoRA) [20], explore which weight matrices in transformers are suitable for the application of LoRA, comprehensive theoretical analyses of these phenomena are still limited. This includes aspects such as selecting appropriate weight types for fine-tuning and optimizing learning rate settings. Reflecting on the attention equation itself (Section 2): (1) In linear algebra, two matrices multiplied without an intermediate activation can be equivalent to a single matrix. Some studies [33, 4] often treat $\mathbf{W}_{q}$ and $\mathbf{W}_{k}$ as a single unit ($\mathbf{W}_{qk}=\mathbf{W}_{q}\mathbf{W}_{k}^{T}$), however, the benefits of fine-tuning $\mathbf{W}_{q}\&\mathbf{W}_{v}$ alone have yet to be further clarified. (2) Considering the scenario where the values of $\mathbf{W}_{q}$, $\mathbf{W}_{k}$, and $\mathbf{W}_{v}$ approach zero, the gradients of $\mathbf{W}_{q}\&\mathbf{W}_{k}$ tend to diminish towards zero. In contrast, the gradient of $\mathbf{W}_{v}$ remains non-zero due to the influence of softmax normalization. Driven by the above motivations, this paper delves into the issue from the following two perspectives.
@@
$\bullet$ Generalization: advantages of fine-tuning $\mathbf{W}_{q}\&\mathbf{W}_{v}$ over $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$ together. We perform a thorough theoretical analysis to demonstrate the advantages. To be more specific, we employ information-theoretic approaches [45, 34, 43, 52] to establish the generalization bounds of fine-tuning pre-trained models with attention mechanism (See Theorem 1 for details). This indicates that fine-tuning $\mathbf{W}_{q}\&\mathbf{W}_{v}$ instead of $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$ reduces the number of parameters, while improving generalization bounds and potentially providing memory benefits.
@@
$\bullet$ Optimization: convergence analysis of attention mechanism with varying learning rate settings. To further investigate the aforementioned phenomena, we examine the optimization process of the attention mechanism. First, we discuss the learning dynamics in transformers in Case 1, suggesting that $\mathbf{W}_{v}$ may experience instances of inefficient learning during downstream task fine-tuning. This naturally leads to the hypothesis that accelerating the learning of $\mathbf{W}_{v}$ in the early stages could potentially induce $\mathbf{W}_{k}$ and $\mathbf{W}_{q}$ to begin learning earlier. Additionally, by using scaling arguments for large width-$n$ networks [47, 12], we illustrate (Theorem 2) that the feature learning of attention mechanism is efficient when the learning rate for $\mathbf{W}_{v}$ should be generally much larger than that of $\mathbf{W}_{q}\&\mathbf{W}_{k}$ in fine-tuning.
@@
Building on our experimental and theoretical insights, one can develop new algorithms to improve the effectiveness (e.g., storage, and time) of fine-tuning. Experimental results for our strategy (in Section 5) on benchmark datasets [44] and open source pre-trained models [28, 2] verify that the method can visibly influence fine-tuning efficiency. We do not make direct comparisons with various parameter-efficient fine-tuning methods, as our strategy is primarily intended to demonstrate how theoretical analysis can effectively guide experimental procedures.
A summary of the main theoretical analyses. According to the traditional statistical learning viewpoint, performance can be defined by the sum of optimization error and generalization error. Our theoretical analyses in Sections 3 and 4 correspond to generalization and optimization, respectively. In Section 3 (generalization, storage-friendly), we give Theorem 1 (information-theoretic generalization bounds), showing that with the same $r$ value, fine-tuning $\mathbf{W}_{q}\&\mathbf{W}_{v}$ consistently achieves results comparable to or even surpassing those of fine-tuning $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$. This reduces the number of parameters for the same $r$, while improving generalization bounds and potentially providing memory benefits. In Section 4 (optimization, time-friendly), we discuss the learning dynamics in fine-tuning attention mechanism, and we illustrate (Theorem 2) that the feature learning of attention mechanism is efficient when the learning rate for $\mathbf{W}_{v}$ should be generally much larger than that of $\mathbf{W}_{q}\&\mathbf{W}_{k}$ in fine-tuning. Building on our experimental and theoretical insights, one can develop new algorithms to improve the effectiveness (e.g., storage, and time) of fine-tuning.
@@ 2 Preliminaries and Background
Self-attention. Given a sequence of $m$ vectors $\mathbf{C}\in\mathbb{R}^{m\times d_{in}}$ over which we would like to perform attention and a query vector $\mathbf{x}\in\mathbb{R}^{d_{in}}$, that is, the input is [$\mathbf{C},\mathbf{x}$] $\in\mathbb{R}^{(m+1)\times d_{in}}$. Conventional attention can be expressed as [^4] For simplicity, we focus on the last vector of input in a single-head self-attention. Our analysis is readily generalizable to multi-head self-attention.:
$\text{Attn}(\mathbf{x}\mathbf{W}_{q},\mathbf{C}\mathbf{W}_{k},\mathbf{C}\mathbf{W}_{v})$
$=\text{softmax}\left(\frac{\mathbf{x}\mathbf{W}_{q}\mathbf{W}_{k}^{T}\mathbf{C}^{T}}{\sqrt{d_{out}}}\right)\mathbf{CW}_{v},$ (1)
@@
A unified framework for parameter-efficient fine-tuning. Building on He et al. [15], we consider a unified framework that establishes connections among various PEFT methods. Specifically, we reinterpret these methods as modifications applied to specific hidden states within pre-trained models, the composition function can be written as:
@@
where $l_{1},l_{2}$ are coefficients, $\mathbf{h}$ is denoted as the hidden representation to be directly modified and $\Delta\mathbf{h}$ is a modification vector. Moreover, $\mathbf{h}$ and $\mathbf{x}$ can represent the attention output and input respectively. Here, we present two special cases:
@@
$\mathbf{h}\leftarrow(1-\alpha(\mathbf{x})\mathbf{h}+\alpha(\mathbf{x})\Delta\mathbf{h},$
$\Delta\mathbf{h}:=\text{softmax}(\mathbf{xW}_{q}\mathbf{P}_{k}^{T})\mathbf{P}_{v}\triangleq\text{softmax}(\mathbf{xA})\mathbf{B},$ (4)
where $\alpha(\mathbf{x})=\frac{\sum_{i}\text{exp}(\mathbf{xW}_{q}\mathbf{P}_{k}^{T})_{i}}{\sum_{i}\text{exp}(\mathbf{xW}_{q}\mathbf{P}_{k}^{T})_{i}+\sum_{j}\text{exp}(\mathbf{xW}_{q}\mathbf{W}_{k}^{T}\mathbf{C}^{T})_{j}}$ is a scalar that represents the sum of normalized attention weights on the prefixes. We derive a detailed equivalent form of Prefix tuning to connect it with LoRA in Appendix A.1.
@@
By defining $\mathbf{A}=\mathbf{W}_{q}\mathbf{P}_{k}^{T},\mathbf{B}=\mathbf{P}_{v}$ in Eq.(4), we can establish a connection with LoRA in Eq.(3). Notably, if we replace the softmax attention with linear attention here, the two are equivalent to some extent. Intuitively, in the attention mechanism, $\mathbf{A}\ (\mathbf{W}_{q}\mathbf{P}_{k}^{T})$ is responsible for generating attention scores, while $\mathbf{B}\ (\mathbf{P}_{v})$ utilizes these attention scores to produce the target content. Therefore, during fine-tuning, query, key, and value are likely to exhibit varying degrees of importance. This may also provide theoretical insights for recent works [52, 11], which empirically observed an asymmetry where the project-down matrix $\mathbf{A}$ is responsible for extracting features from the input, while the project-up matrix $\mathbf{B}$ utilizes these features to generate the desired output in LoRA.
$\Theta$ Notation. In our convergence analysis, we use standard asymptotic notation to describe behavior as the width $n$ grows, following conventions in [47, 12]. Given sequences $c_{n}\in\mathbb{R}$ and $d_{n}\in\mathbb{R}^{+}$, we write $c_{n}=O(d_{n})$ and $c_{n}=\Omega(d_{n})$ to mean $c_{n}<\kappa d_{n}$ or $c_{n}>\kappa d_{n}$, respectively, for some constant $\kappa>0$. We denote $c_{n}=\Theta(d_{n})$ when both $c_{n}=O(d_{n})$ and $c_{n}=\Omega(d_{n})$ hold, implying that $c_{n}$ and $d_{n}$ grow at comparable rates. For vector sequences $c_{n}=(c_{n}^{i})_{1\leq i\leq k}\in\mathbb{R}^{k}$ (for some $k>0$), we write $c_{n}=O(d_{n})$ when $c_{n}^{i}=O(d_{n}^{i})$ for all $i\in[k]$, and analogous notation applies for other asymptotic bounds. Finally, when the sequence $c_{n}$ is a vector of random variables, convergence is understood to refer to convergence in the second moment (i.e., $L_{2}$ norm).
@@ 3 Advantages and Generalization Analysis
In this section, we show our first interesting observation (Unequal Importance of Attention Matrices) in fine-tuning the attention mechanism and the storage benefits of fine-tuning only $\mathbf{W}_{q}$ and $\mathbf{W}_{v}$ (Section 3.1). Next, we give mutual information based generalization bounds of fine-tuning only $\mathbf{W}_{q}$ and $\mathbf{W}_{v}$ (Section 3.2), providing a better generalization error.
3.1 Empirical Advantages
To explore the Unequal Importance of Attention Matrices, we focus our study on adapting only the attention weights for downstream tasks, while freezing the other modules to ensure simplicity and parameter efficiency. Furthermore, we investigate the impact of adapting different types of attention weight matrices in a Transformer, as outlined below. We present our empirical results using LoRA to fine-tune a set of language models (Roberta-base [28] and Llama3.1-8b [2]) across various benchmarks [44]. Further details on the experimental setup and additional empirical results are in Appendix B.1.
Table 1 provides a detailed comparison of the impact of fine-tuning different weight matrices ($\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$) across various rank values $r$ and weight update strategies in LoRA fine-tuning on tasks like SST2, QNLI, QQP, and MNLI. We can see a clear trend where solely updating the $\mathbf{W}_{v}$ matrix outperforms just learning the $\mathbf{W}_{q},\mathbf{W}_{k}$ matrix. Interestingly, the combination of fine-tuning both $\mathbf{W}_{q}$ and $\mathbf{W}_{v}$ often leads to performance that matches or even exceeds that achieved by fine-tuning all three matrices $\mathbf{W}_{q},\mathbf{W}_{k},$ and $\mathbf{W}_{v}$. This pattern is consistently observed across various tasks and rank values, further emphasizing the importance of these two matrices over $\mathbf{W}_{k}$ during fine-tuning.